Menlo Company distributes a single product. The companys sales and expenses for last month follow: Per Unit Total Sales VariReq 2 Req 3A Req 1 Req 3B Req 4 Req 5 What is the monthly break-even point in unit sales and in dollar sales? Break-even poinReq 1 Req 2 Req 3A Without resorting to computations, what is th Total contribution marginReq 1 Req 2 Req 3A Req 3B Req 4 Req 5 How many units would have to be sold each month to attain a target profit of $38,400? UReq 3A Req 1 Req 2 Req 3B Req 4 Req 5 Verify your answer by preparing a contribution format income statement at the target saReq 3A Req 1 Req 2 Req 3B Req 4 Req 5 Refer to the original data. Compute the companyss margin of safety in both dollar andReq 5 Req 1 Req 2 Req 3A Req 3B Req 4 What is the companys CM ratio? If sales increase by $77,000 per month and there is no
Answer #1

1.Monthly break even point in unit sales = Fixed costs/Contribution Margin per unit

= 75,000/6

= 12,500 units

In Dollar Sales = Fixed costs/Contribution Margin Ratio

Contribution Margin Ratio = Contribution Margin/Sales

= 6/20 = 30%

Hence, break even sales = 75,000/30%

= $250,000

2.Total contribution margin at break even point = fixed costs

= $75,000

3-a. Target Profit = $38,400

Add: Fixed costs = $75,000

Desired Contribution Margin = $113,400

Units required to be sold = 113,400/6

= 18,900 units

3-b. Contribution Format Income Statement

Total

Per Unit

Sales

378,000

20

Variable Expenses

264,600

14

Contribution Margin

113,400

6

Fixed Expenses

75,000

Net Operating Income

38,400

4.Margin of Safety in dollars = Sales – Break even sales

= 308,000-250,000

= $58,000

In % = 58,000/308,000

= 18.83%

5.CM Ratio = 6/20

= 30%

Monthly net operating income will increase by the amount of increase in contribution margin

= 77,000*30%

= $23,100